OpenStudy (anonymous):

g(x)=|3x+3|+3 find g(-4)

ganeshie8 (ganeshie8):

plugin x = -4 in the given equaiton

ganeshie8 (ganeshie8):

g(x)=|3x+3|+3 g(-4) = |3(-4) + 3| + 3 = |-12+3| + 3 = ?

OpenStudy (gorv):

and remember mod ealwys give positive value

OpenStudy (gorv):

like if |-2|=2